In a box there are 4 blue and 8 white marbles. If 2 marables are selected at random, what is the probability to see


a, marbles of the same color?


b, marbles of different color?


c, both blue marbles ?


d, both white marbles?


1
Expert's answer
2022-01-31T16:36:25-0500

There are "4+8=12" marbles.


"S=\{BB, BW, WB, WW\}"

a)


"P(2\ same \ color)=P(BB)+P(WW)"

"=\dfrac{4}{12}(\dfrac{4-1}{12-1})+\dfrac{8}{12}(\dfrac{8-1}{12-1})=\dfrac{17}{33}"

b)


"P(2\ different \ color)=P(BW)+P(WB)"

"=\dfrac{4}{12}(\dfrac{8}{12-1})+\dfrac{8}{12}(\dfrac{4}{12-1})=\dfrac{16}{33}"

"=1-P(2\ same \ color)"

c)


"P(BB)=\dfrac{4}{12}(\dfrac{4-1}{12-1})=\dfrac{1}{11}"

d)


"P(WW)=\dfrac{8}{12}(\dfrac{8-1}{12-1})=\dfrac{14}{33}"